OpenStudy (anonymous):

A hyperbola has vertices (+-5,0) and one focus (6,0) what is the standard-form equation of the hyperbola?

OpenStudy (rsadhvika):

c^2 = a^2 + b^2

OpenStudy (rsadhvika):

6^2 = 5^2 + b^2 b^2 = 11

OpenStudy (rsadhvika):

\(\large \frac{x^2}{25} - \frac{y^2}{11} = 1\)